CAPM and required return

Beale Manufacturing Company has a beta of 2.5, and Foley Industries has a beta of 0.3. The required return on an index fund that holds the entire stock market is 9%. The risk-free rate of interest is 3.75%. By how much does Beale's required return exceed Foley's required return? Round your answer to two decimal places.
Answer #1

required return=risk-free rate +Beta*(market rate- risk-free rate)

required return for:

Beale Manufacturing Company=3.75+(9-3.75)*2.5=16.875%

Foley Industries=3.75+(9-3.75)*0.3=5.325%

Hence difference=16.875%-5.325%

=11.55%